OpenStudy (anonymous):

Georgia has only dimes and quarters in a bag. She has a total of 18 coins that are worth $3. How many more dimes than quarters does she have?

OpenStudy (anonymous):

let D = the number of dimes let Q = number of quarters we know from the problem that the number of coins is 18 so we can safely say that D + Q = 18

OpenStudy (anonymous):

yeah.... i got that far then i did .25q + .10d = 3.00 and well.... it doesn't work.

OpenStudy (anonymous):

try setting D = 18 - Q and then plug D into ^ equation

OpenStudy (anonymous):

i did, i got some odd equation and since it's asking how many, an eqation of any sort cannot be an answer

OpenStudy (anonymous):

lets work through it step by step if I let D = 18-Q and then plug in we get .25Q + .10(18-Q) = 3.00

OpenStudy (anonymous):

then we distribute the .10 so .25Q - 1.80 - .10Q = 3.00

OpenStudy (anonymous):

we can then combine like terms .15Q -1.80 = 3.00

OpenStudy (anonymous):

.15q - 4.8*

OpenStudy (anonymous):

=*

OpenStudy (anonymous):

no wait i made a mistake it should be .25Q + 1.80 -.10Q = 3.00 so you were right .15Q = 1.2

OpenStudy (anonymous):

and once you find a value for Q you can plug that into D + Q = 18 to find D

OpenStudy (anonymous):

and then take the difference D - Q to find how many more dimes than quarters there are

OpenStudy (anonymous):

myeh -.- i did the same thing and got it wrong, with you i got it right myeh myeh myeh! jubjub! i hate math >.> anyway thanks

OpenStudy (anonymous):

np i made a mistake too